OpenStudy (anonymous):

Find the value of the x and y determinants in the following system: -3x+2y=-6 4x-3y=4

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

Equation 1 * 3 = -9x + 6y = -18 Equation 2 * 2 = 8x - 6y = 8 then add them together to get -x + 0y = -10 divde by -1 x = 10

OpenStudy (anonymous):

I'm so confused.... What am I supposed to do? How did you arrive at that?

OpenStudy (anonymous):

Wait.... Is this the elimination method?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

okay, I think I see what you did...

OpenStudy (anonymous):

what about the determinant part?

OpenStudy (anonymous):

isn't that bª-4ac?

OpenStudy (anonymous):

I think it just wants the solution of the equation.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Because we know x = 10 we can plug it into the first equation of -3x+2y=-6 -3 * 10 +2y = -6 -30 + 2y = -6 2y = 24 y = 12

OpenStudy (anonymous):

okay, that makes sense!

OpenStudy (anonymous):

Do you think you could help me with a few more questions? I can post them as new ones

OpenStudy (anonymous):

Sure